Find a linear approximation to 4.1. Please enter your answer below and upload your solution
Answer #1

Solution: V4.1 L(x) = f(a) + f(a)(x-a) f(x) = VX f(x) = z Va L(x) = V4+ztax – 4) taking a = 4 => L(4.1) = 2 + 2(4.1 – 4) => L

please comment for queries.

thanks.